Calculation of individual costs and WACC   Dillon Labs has asked its financial manager to measure the cost of each specific type of capital as well as the weighted average cost of capital. The weighted average cost is to be measured by using the following​ weights:35​%Long-term debt, 15​% preferred​ stock, and 50​% common stock equity​ (retained earnings, new common​ stock, or​ both). The​ firm's tax rate is 22​%.

.Debt The firm can sell for ​$1020 a 16​-year, ​$1,0001,000​-par-value bond paying annual interest at a 11.00​% coupon rate. A flotation cost of 33​% of the par value is required.

Preferred stock 9.00​% ​(annual dividend) preferred stock having a par value of $100 can be sold for ​$94. An additional fee of ​$55 per share must be paid to the underwriters.

Common stock  The​ firm's common stock is currently selling for ​$80 per share. The stock has paid a dividend that has gradually increased for many​ years, rising from ​$2.00 ten years ago to the $3.58 dividend​ payment,Upper D 0D0​, that the company just recently made. If the company wants to issue new new common​ stock, it will sell them

​$1.50 below the current market price to attract​ investors and the company will pay ​$4.00 per share in flotation costs.  

a.  Calculate the​ after-tax cost of debt.

b.  Calculate the cost of preferred stock.

c.  Calculate the cost of common stock​ (both retained earnings and new common​ stock).
Answer #1

cost of Debt = 6 + Mr P. Cool Map L Whing e= coupon payment NE Maturity haul Po Not proceeds. 1020-330

& Costo Debt 2013 110 t 1.000 76902 | - 1 16 , X 100 jooo 690 121.375 X100 845 -ot after to cast 15.31.01 10:22) t of Debit a

TALO. Pois 20-115-12-1 | Cost of common stoas e Divity AT ASOSPOL POTENT DE - Pos 80-115-pipeline ) - 74.5 growth oute = /3.5

wiegnted awedase cost of capital E (0.35Xll. ghy+ C0.15X23-087.)+20.5*11.0911 | ! 13:1860%

I hope my efforts will be fruitful to you...?